# Analysis of a multi-target linear shrinkage covariance estimator

### v1

@@ 3 Oracle and proposed estimator
$c_{0}$ $=\min\left((\hat{c}_{0})_{+},1\right),\text{ with }\hat{c}_{0}=\frac{1}{\det(A_{n})}\left(\lVert S_{n}\rVert_{n}^{2}-\hat{\mathbb{V}}(S_{n})-\sum_{k=1}^{N_{n}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}-\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)\right)\right),$ (14)
$c_{i}$ $=(1-c_{0})\langle S_{n},T_{n}^{(i)}\rangle_{n},$
The MTSE is defined as: $S_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}S_{n}+(1-c_{0})\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}\right).$
The truncation of $c_{0}$ between 00 and $1$ theoretically does not change the asymptotic behavior of the estimator. Geometrically, it ensures that $S_{n}^{*}$ is a convex combination between $S_{n}$ and the matrix $\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}$. Empirically, it improves the robustness of the estimator, particularly in small sample size.
Immediately from Theorem 1, we have the following result.
Corollary 1.
Under assumptions 1 to 4, we have $(\hat{c}_{1}-c_{1}^{*})\det(A_{n})\underset{q.m}{\longrightarrow}0$.
Theorem 2 (Loss convergence).
The idea of the proof is to prove the convergence on the estimators without the projection step onto the cone of positive matrices, and use the Lipschitz property of projection to conclude. The initial convergence relies on the Lemma A.1 in [13]. The check of the hypotheses for the use of the Lemma comes essentially from the previous results we proved.
4 Experimental results
This section illustrates the practical consequences of the choice of targets on the estimation, the robustness of the estimator to heavy tails, and an application to real non-stationary time series. The experiments were computed on a Lenovo laptop, Intel i5, 32Go Ram. The performance is measured with the Percentage Relative Improvement in Average Loss (PRIAL):
4.1 Target set impact: Usefulness of a target
We fix the dimension at $p=50$, and $n=25$, and we study the impact of the number of targets $N_{n}$. The population covariance $\Sigma$ is designed to clearly exhibit a choice of useful targets, and compare the role of targets that give information on the population covariance - i.e. that improve significantly the oracle PRIAL-, to those which give less. $\Sigma$ is built as follows: this is a block diagonal matrix composed of 10 block matrices $(B_{i})_{i=1}^{10}$ of shape $(5,5)$. For each $i\in\llbracket 1,10\rrbracket$, $B_{i}$ is drawn from a Wishart distribution with $5$ degrees of freedom and $\sqrt{11-i}I_{5}$ as scale matrix.
We consider two sets of targets: one set of “aligned” targets $(T_{1})^{(i)}$, and one set of “misaligned” ones $(T_{2})^{(i)}$. The first one is built as following: the first target $T_{1}^{(1)}=I$, to match Ledoit-Wolf STSE [13], and for $i\in\llbracket 2,10\rrbracket$, the target $T_{1}^{(i)}$ is block-diagonal; where $\Sigma$ has the block $B_{i}$, $T_{1}^{(i)}$ has an Identity block; elsewhere $T_{1}^{(i)}$ has null values.
The set of target is orthogonalized with the Gram-Schmidt algorithm.
For the second set of targets $(T_{2})^{(i)}$, we break the perfect alignment between $T_{1}^{(i)}$ and the block $B_{i}$. For that, we apply a permutation $\sigma$ on both rows and columns of $T_{1}^{(i)}$, where $\sigma$ is defined as:
We consider the PRIAL of the MTSE (MTSE_1 and MTSE_2 respectively for target set 1 and 2), the oracle ones, and the Ledoit-Wolf STSE (LW), in function of the number $k$ of targets used, between $k=1$, which is equivalent to the STSE, and $k=10$ where the sets of targets $(T_{1}^{(i)})_{i=1}^{k}$ and $(T_{2}^{(i)})_{i=1}^{k}$ are used completely. The results, with $50000$ draws, are shown in Figure 2. The maximum standard deviation is $4\times 10^{-4}$, error bars were removed.
Figure 1: (left) PRIAL of the MTSE, and the oracle, with target sets 1 and 2, and the Ledoit-Wolf STSE as reference, in function of the number of targets in the experiment of target usefulness.
Figure 2: (right) PRIAL difference between the oracle MTSE and Ledoit Wolf linear shrinkage one (MTSE_o - LW) in orange, and estimator MTSE and Ledoit Wolf linear shrinkage one (MTSE - LW) in blue, in function of the number of targets in the setup of the experiment of target set impact.
It confirms the intuition about what is a useful target: the “aligned” set $(T_{1})^{(i)}$ presents a significant better PRIAL than the second misaligned one $(T_{2})^{(i)}$. This comprehension of what makes a target useful is of utter importance to design the target set in a practical use case. The second conclusion is that even if the misaligned targets are not optimal, they still present a better PRIAL than the STSE using only one target.
4.2 Target set impact: Adding useless targets to a good target set
The first target is the identity matrix $I$, to match with Ledoit-Wolf STSE. For $i\in\llbracket 2,10\rrbracket$, $T^{(i)}$ is designed according to Equation 17. For target $i\geq 11$, the target is drawn from a Wishart distribution with $p$ degrees of freedom and $I_{p}$ as scale matrix. The latter are supposed to be “useless” regarding the structure of $\Sigma$. The set of targets is then orthogonalized with the Gram-Schmidt algorithm.
[…]

### v2

@@ C.2 Proof of Lemma 1
$\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T\rangle_{n}^{4}\right]=$ $\frac{1}{n^{4}}\sum_{k_{1},k_{2},k_{3},k_{4}}\mathbb{E}\left[\prod_{m\in\{k_{1},k_{2},k_{3},k_{4}\}}\left\langle X_{m}^{n}\left(X_{m}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\right]$ (22)
$\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T\rangle_{n}^{4}\right]\leq$ $\frac{M^{\prime}n(n-1)}{n^{4}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\rangle_{n}^{2}\right]^{2}$
$+\frac{n}{n^{4}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\rangle_{n}^{4}\right].$
Moreover, we have:
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{4}\right]$ $\leq\mathbb{E}\left[\left(\sum_{k}\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{2}\right)^{2}\right]$ (23)
$\leq\mathbb{E}\left[\left\lVert X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n}\right\rVert_{n}^{4}\right]$
$\leq\mathbb{E}\left[\left\lVert X_{1}^{n}\left(X_{1}^{n}\right)^{T}\right\rVert_{n}^{4}\right]$
$=\frac{1}{p_{n}^{2}}\mathbb{E}\left[\left(\sum_{i}y_{i1}^{2}\right)^{4}\right]$
$=p_{n}^{2}\mathbb{E}\left[\left(\frac{1}{p_{n}}\sum_{i}y_{i1}^{2}\right)^{4}\right]$
$\leq p_{n}^{2}\frac{1}{p_{n}}\sum_{i}\mathbb{E}\left[y_{i1}^{8}\right]$
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{4}\right]$ $\leq p_{n}^{2}K_{2}.$
And,
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right]^{2}$ $\leq\max_{i}\left(\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(i)}\rangle_{n}^{2}\right]\right)\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right]$ (24)
$\leq\max_{i}\left(\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(i)}\rangle_{n}^{2}\right]\right)\mathbb{E}\left[\lVert X_{1}^{n}\left(X_{1}^{n}\right)^{T}\rVert_{n}^{2}\right]$
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right]^{2}$ $\leq\max_{i}\left(\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T_{n}^{(i)}\rangle_{n}^{2}\right]\right)p_{n}\sqrt{K_{2}}.$
For any of the targets $T$, we have with Assumption 2: $\mathbb{E}\left[\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T},T\rangle_{n}^{2}\right]\leq\sqrt{K_{2}}$. Finally, using $M=M^{\prime}+K_{1}$, we have:
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{4}\right]\leq\frac{M^{\prime}+K_{1}}{n}K_{1}K_{2}=\frac{M}{n}K_{1}K_{2}.$ (25)
Unkown mean
When the mean is unknown, we have a larger decomposition. We use the same technics as in the known mean setting, with more terms and cross-products to handle:
$\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T\rangle_{n}^{4}\right]=$ (26)
$\frac{1}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\mathbb{E}\left[\prod_{s=1}^{4}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\right]$ Term (1)
$-4\mathbb{E}\left[\prod_{s=1}^{3}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{k_{4}}^{n}\left(X_{k_{4}^{\prime}}^{n}\right)^{T}-\delta_{k_{4}=k_{4}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]$ Term (2)
$+6\mathbb{E}\left[\prod_{s=1}^{2}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\prod_{t=3}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]$ Term (3)
$-4\mathbb{E}\left[\left\langle X_{k_{1}}^{n}\left(X_{k_{1}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\prod_{t=2}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]$ Term (4)
$+\mathbb{E}\left[\prod_{t=1}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right].$ Term (5)
We prove the higher bound of the form $M^{\prime\prime}K_{1}K_{2}/n$ term by term, in absolute value, summing over all targets $(T_{n}^{(k)})_{k=1}^{N_{n}}$. For that, we decompose each term and bound them by expectations we can control.
Term (1)
This term is equal to $\frac{n^{4}}{(n-1)^{4}}\mathbb{E}\left[\langle S_{n}^{(\text{known mean)}}-\Sigma_{n},T\rangle_{n}^{4}\right]$. Supposing $n\geq 4$, we have $\frac{n^{4}}{(n-1)^{4}}\leq 4$, so, with $M_{1}^{\prime\prime}=4M$, we have:
$\sum_{k=1}^{N_{n}}\frac{1}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{s=1}^{4}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\right]\right|\leq\frac{M^{\prime\prime}_{1}}{n}K_{1}K_{2}.$ (27)
Term (2)
For this term, we have the following development, when discarding 00-expectation terms:
$\frac{4}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{s=1}^{3}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{k_{4}}^{n}\left(X_{k_{4}^{\prime}}^{n}\right)^{T}-\delta_{k_{4}=k_{4}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (28)
$\leq$ $\frac{4}{n}\times\left|\text{Term(1)}\right|+\frac{24}{(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\left\langle X_{2}^{n}\left(X_{2}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}\right]\right|$
$\leq$ $\frac{4}{n}\times\left|\text{Term(1)}\right|+\frac{24}{(n-1)^{3}}\sqrt{\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{4}\right]^{3/2}\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]}$
$\leq$ $\frac{4}{n}\times\left|\text{Term(1)}\right|+\frac{6}{(n-1)^{3}}\left(3\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{4}\right]+\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]^{2}\right).$
At this point, we recall that we have to sum over each target $(T_{n}^{(k)})_{k=1}^{N_{n}}$ and then bound by a term of the form $M^{\prime\prime}_{2}K_{1}K_{2}/n$ for some $M_{2}^{\prime\prime}\in\mathbb{R}$. Term(1) is obviously controlled by the previous part. For the rest, we use the two following inequalities:
  • from the proof in the known mean setting, we have that:
    $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T_{n}^{(k)}\right\rangle_{n}^{4}\right]\leq p_{n}^{2}K_{2},$ (29)
  • and:
    $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right]^{2}$ (30)
    $\leq\max_{i}\left(\mathbb{E}\left[\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(i)}\rangle_{n}^{2}\right]\right)\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right]$
    $\leq\max_{i}\left(\mathbb{E}\left[\frac{1}{p_{n}^{2}}\left(X_{2}^{n}\right)^{T}T_{n}^{(i)}X_{1}^{n}\left(X_{1}^{n}\right)^{T}T_{n}^{(i)}X_{2}^{n}\right]\right)\mathbb{E}\left[\lVert X_{1}^{n}\left(X_{2}^{n}\right)^{T}\rVert_{n}^{2}\right]$
    $\leq\max_{i}\left(\frac{1}{p_{n}}\left\langle\Sigma_{n}T_{n}^{(i)},T_{n}^{(i)}\Sigma_{n}\right\rangle_{n}\right)p_{n}\sqrt{K_{2}}$
    $\leq\max_{i}\left(\frac{1}{p_{n}}\left\lVert\Sigma_{n}T_{n}^{(i)}\right\rVert_{n}^{2}\right)p_{n}\sqrt{K_{2}}$
    $\leq\left\lVert\Sigma_{n}\right\rVert_{n}^{2}p_{n}\sqrt{K_{2}}$
    $\leq p_{n}K_{2}.$
Supposing $n\geq 4$, with $M_{2}^{\prime\prime}=M_{1}^{\prime\prime}+54K_{1}$, we then have:
$\sum_{k=1}^{N_{n}}\frac{4}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{s=1}^{3}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{k_{4}}^{n}\left(X_{k_{4}^{\prime}}^{n}\right)^{T}-\delta_{k_{4}=k_{4}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (31)
$\leq\frac{M^{\prime\prime}_{2}}{n}K_{1}K_{2}.$
Term (3)
Similarly for this term, we obtain:
$\frac{6}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{s=1}^{2}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\prod_{t=3}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (32)
[…]
$+\frac{12(n-2)}{n(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\right]\right|$
$+\frac{48(n-2)}{n(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{2}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}\right]\right|$
$+\frac{24}{n(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{2}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]\right|$
$+\frac{24}{n(n-1)^{3}}\left||\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]\right|$
$\leq$ $\frac{6(2n-1)}{n^{2}}\times\left|\text{Term(2)}\right|+\frac{24}{n(n-1)^{2}}\left(\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{4}\right]+\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{4}\right]\right)$
$+\frac{6(n-2)}{n(n-1)^{3}}\left(\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\right]^{2}+\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]^{2}\right).$
We use the two following inequalities to finish the work, combined with the two ones we used for Term (2):
  • from the proof in the known mean setting, we have that:
    $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T_{n}^{(k)}\right\rangle_{n}^{2}\right]^{2}\leq p_{n}K_{2},$ (33)
  • and:
    $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{4}\right]$ $\leq\mathbb{E}\left[\left(\sum_{k=1}^{N_{n}}\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{2}\right)^{2}\right]$ (34)
    $\leq\mathbb{E}\left[\left\lVert X_{1}^{n}\left(X_{2}^{n}\right)^{T}\right\rVert_{n}^{4}\right]$
    $=\frac{1}{p_{n}^{2}}\mathbb{E}\left[\left(\sum_{i}y_{i1}y_{i2}\right)^{4}\right]$
    $\leq p_{n}^{2}\frac{1}{p_{n}}\sum_{i}\mathbb{E}\left[y_{i1}^{4}y_{i2}^{4}\right]$
    $\leq p_{n}^{2}\frac{1}{p_{n}}\sum_{i}\mathbb{E}\left[y_{i1}^{8}\right]$
    $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T_{n}^{(k)}\rangle_{n}^{4}\right]$ $\leq p_{n}^{2}K_{2}.$
Supposing $n\geq 4$, with $M_{3}^{\prime\prime}=3M_{2}^{\prime\prime}+86K_{1}+1$, we then have:
$\sum_{k=1}^{N_{n}}\frac{6}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{s=1}^{2}\left\langle X_{k_{s}}^{n}\left(X_{k_{s}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\prod_{t=3}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (35)
$\leq\frac{M^{\prime\prime}_{3}}{n}K_{1}K_{2}.$
Term (4)
For that one, we have:
$\frac{4}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\left\langle X_{k_{1}}^{n}\left(X_{k_{1}}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\prod_{t=2}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (36)
[…]
$+\frac{96(n-2)}{n^{2}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{3}^{n}\left(X_{1}^{n}\right)^{T},T\right\rangle_{n}\right]\right|$
$+\frac{96(n-2)}{n^{2}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]\right|$
$+\frac{24}{n^{2}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{3}\right]\right|$
[…]
$\frac{48n-88}{n^{2}(n-1)^{3}}\left(3\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{4}\right]+\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{1}^{n}\right)^{T}-\Sigma_{n},T\right\rangle_{n}^{2}\right]^{2}\right)$
Supposing $n\geq 4$, with $M_{4}^{\prime\prime}=M_{3}^{\prime\prime}+47K_{1}+4$, we then have:
$\sum_{k=1}^{N_{n}}\frac{4}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\left\langle X_{k_{1}}^{n}\left(X_{k_{1}}^{n}\right)^{T}-\Sigma_{n},T_{k}^{(n)}\right\rangle_{n}\prod_{t=2}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T_{k}^{(n)}\right\rangle_{n}\right]\right|$ (37)
$\leq\frac{M^{\prime\prime}_{4}}{n}K_{1}K_{2}.$
Term (5)
Following the same idea, we have for the last term:
$\frac{1}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{t=1}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T\right\rangle_{n}\right]\right|$ (38)
$\leq$ $\frac{n^{3}+n^{2}-4}{n^{4}}\times\left|\text{Term(4)}\right|+\frac{8}{n^{3}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{4}\right]\right|$
$+\frac{24(n-2)}{n^{3}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\left\langle X_{1}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]\right|$
$+\frac{12(n-2)(n-3)}{n^{3}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\left\langle X_{3}^{n}\left(X_{4}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]\right|$
$+\frac{48(n-2)(n-3)}{n^{3}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{3}^{n}\left(X_{1}^{n}\right)^{T},T\right\rangle_{n}\right]\right|$
$+\frac{48(n-2)(n-3)}{n^{3}(n-1)^{3}}\left|\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{2}^{n}\left(X_{3}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{3}^{n}\left(X_{4}^{n}\right)^{T},T\right\rangle_{n}\left\langle X_{4}^{n}\left(X_{1}^{n}\right)^{T},T\right\rangle_{n}\right]\right|$
$\leq$ $\frac{n^{3}+n^{2}-4}{n^{4}}\times\left|\text{Term(4)}\right|+\frac{72n+8}{n^{3}(n-1)^{3}}\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{4}\right]$
$+\frac{60(n-2)(n-3)}{n^{3}(n-1)^{3}}\mathbb{E}\left[\left\langle X_{1}^{n}\left(X_{2}^{n}\right)^{T},T\right\rangle_{n}^{2}\right]^{2}.$
Supposing $n\geq 4$, with $M_{5}^{\prime\prime}=M_{4}^{\prime\prime}+11K_{1}+2$, we then have:
$\sum_{k=1}^{N_{n}}\frac{1}{n^{4}(n-1)^{4}}\sum_{\begin{subarray}{c}k_{1},k_{2},k_{3},k_{4}\\ k_{1}^{\prime},k_{2}^{\prime},k_{3}^{\prime},k_{4}^{\prime}\end{subarray}}\left|\mathbb{E}\left[\prod_{t=1}^{4}\left\langle X_{k_{t}}^{n}\left(X_{k_{t}^{\prime}}^{n}\right)^{T}-\delta_{k_{t}=k_{t}^{\prime}}\Sigma_{n},T_{k}^{(n)}\right\rangle_{n}\right]\right|\leq\frac{M^{\prime\prime}_{5}}{n}K_{1}K_{2}.$ (39)
Finally, combing the 5 terms, we have with $M^{\prime\prime}=\sum_{i=1}^{5}M_{i}^{\prime\prime}$:
$\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{4}\right]\leq\frac{M^{\prime\prime}}{n}K_{1}K_{2},$ (40)
which concludes the proof for the unknown mean.
C.3 Proof of Theorem 1
Known mean
When the mean is known, $\hat{\mathbb{V}}(S_{n})$ is trivially unbiased when developping. $\mathbb{E}\left[\left(\hat{\mathbb{V}}(S_{n})-\mathbb{V}(S_{n})\right)^{2}\right]\longrightarrow 0$ is a consequence of Lemma 3.4 in [15] about the quadratic mean convergence of $\bar{b}_{n}^{2}-\beta_{n}^{2}=\frac{n-1}{n}\hat{\mathbb{V}}(S_{n})-\mathbb{V}(S_{n})$.
$\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ is trivially unbiased when developing. Let’s prove that $\mathbb{E}\left[\left(\sum_{i}\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})-\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n})\right)^{2}\right]\longrightarrow 0$.
[…]